Grade 11Physical Chemistry

The specific charge of a proton is 9.6 x 10^7 C kg^-1, then for an alpha particles it will be: (1) 2.4 x 10^7 C kg^-1 (2) 4.8 x 10^7 C kg^-1 (3) 19.2 x 10^7 C kg^-1 (4) 38.4 x 10^7 c kg^-1Explain???

Nitin Kumar
9 Years agoGrade 11

4 Answers

Vikas TU
9 Years ago
Alpha particle is nothing but He atom.
And He atom has 2 as its atomic number.
And therfore would have 2 protons.
Hence, the total charge would be = > 2 x 9.6 x 10^7 C kg^-1 = > 19.2 x 10^7 kg^-1
Ashish
7 Years ago
4.8×10^7 C kg—1. It's the right answer.    Alpha particle has 2 protons.    9.6×10^7/4 . It is the right answer. 
Soumya
7 Years ago
For proton e/m = 9.6 x 10^7 C kg^-1
For alpha particle = (2/4)e/m = (1/2)e/m
= 0.5 * 9.6 x 10^7 C kg^-1
=4.8 x 10^7 C kg^
Subhadeep
7 Years ago
Given that, specific proton has charge 9.6×10^7c/kg. Alpha particles are nothing but helium(he) atoms." He" has 2 protons (atomic number)  . so Alpha particles has : 2×9.6×10^7 c/kg i.e =19.2×10^7c/kg